OpenStudy (anonymous):

3x^4+9x^2+5x --------------- 3x

sam (.sam.):

split and solve 3x^4/3x+9x^2/3x+5x/3x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x^3+3x+5/3

OpenStudy (anonymous):

start with 3x^4!

sam (.sam.):

\[\frac{3x^4}{3x}+\frac{9x^2}{3x}+\frac{5x}{3x}\]

OpenStudy (anonymous):



OpenStudy (anonymous):